OpenStudy (anonymous):

A 12-ounce Pepsi contains 54 mg of caffeine. What is the average caffeine content per ounce of Pepsi? Round your answer to the nearest tenth, if needed.

OpenStudy (anonymous):

\[\frac{ 12 oz }{ 54 mg }= \frac{ 1 oz }{ xmg }\] 12/12 = 1. So 54/12= 4.5. So in one ounce of pepsi there is 4.5mg of caffeine.